Status Report

ISS On-Orbit Status 12 Jun 2002

By SpaceRef Editor
June 12, 2002
Filed under , ,

All ISS systems continue to function nominally,
except as noted previously or below.   It is Day 189 for
the Expedition 4 crew (187 days on board ISS).

Station day began at 5:53 am EDT (Shuttle day 30 min.
earlier), and the crews awoke to another message of emphatic kudos
for a great EVA-2 day yesterday.  Congratulations were also
extended to Carl Walz and Dan Bursch for breaking the previous U.S.
record for longest space flight last night at 10:19 pm EDT, when they
passed the 188d 4h mark set by Dr. Shannon Lucid on Mir.  This
morning at 2:55 am, Walz also took over the U.S. record for total
space flight time.

Crew activities today centered largely on cleanup from yesterday’s
EVA, crew experience handovers, and MPLM (multi-purpose logistics
module) transfers.  The location map for the final MPLM stowage
layer 3B was uplinked ahead of time, and planned return mass
properties are within latest requirement for the module’s
center-of-gravity location. Transfers today dealt with layer 3B
items, and stowage of return cargo should be finished well ahead of
time.  Return of MPLM Leonardo to the Shuttle cargo bay is
planned for tomorrow afternoon.

At 9:08 am EDT today, Endeavour performed the second automated
one-hour reboost of the station.  Attitude control was handed
over to the Shuttle at 8:27 am, and the "stack" was
maneuvered to the reboost attitude at 8:32 am (LVLH Yaw/Pitch/Roll =

CDR Valery Korzun, with Yuri Onufrienko at his side, completed the
regular daily servicing tasks of SOSh life support systems
maintenance (incl. toilet systems), Lab payloads status check, and
IMS (inventory management system) delta file preparation.

Dan Bursch performed the morning and evening descent checks of the
BPS (biomass production system), while Carl Walz checked the other
active Shuttle middeck payloads.

EXPRESS Rack 1 and SAMS (space acceleration measurement system) are
complete with their activities, and the ground had powered the rack
off by 7:00 am EDT.

The crewmembers performed their daily physical exercise on TVIS
(treadmill) and CEVIS (cycle), reduced to one-hour sessions.

During EVA-2, when the spacewalkers tried to re-install the TDA (TUS
[trailing umbilical system] disconnect actuator) safing bolt of the
Mobile Transporter’s IUA (interface umbilical assembly), they only
got a total of 3.75 turns out of the expected 7 turns.  It
appears that a minimum of 6.25 turns is required to safe the unit.
 Therefore, the device is currently in an undetermined state and
an engineering assessment indicates that the bolt would probably not
be able to stop the cable cutter blade if the device fired.  The
issue with the safing bolt surfaced during EVA-2 of Mission 8A on

During EVA-2 preparations, done on PSA (power supply assembly) power
for keeping the EMU (extravehicular mobility unit, = spacesuit)
batteries charged, the PSA tripped on both EMU 1 and 2 channels, when
the umbilical of one of the EMUs  was disconnected, switching
the suit from local power to battery power.  No current
anomalies were seen; the power-down was caused by a voltage spike
which triggered an overvoltage switch.  Since the EMU battery
had sufficient charge, the crew proceeded with the EVA-2.  The
PSA anomaly has been seen before on EVA-1 (with only one of the two
channels lost), but the reason for it is not understood, and an
investigation is underway.

In a second occurrence, EV1’s EMU battery indicated an increased
discharge rate at the end of the spacewalk.  The battery
(#2022), which had been used without incident on EVA-4 of Mission 8A,
will be returned to Earth on UF-2 for analysis.  It is certain
that the battery voltage problem and the PSA EMU1 and 2 voltage limit
trip are not related.  In light of both issues, the batteries
for EVA-3 are being charged today in the Shuttle airlock; in
addition, it is considered to do tomorrow’s EVA preparations on
battery power and then replace EMU batteries before crew egress from
the Joint Airlock (JAL) with fresh batteries.  Final decision is
being developed this afternoon.

The planned Metox (metal oxide) canister regeneration last night was
not performed, due to concern about a "wet" indication by a
sensor of the JAL’s CCAA (common cabin air assembly) heat exchanger.
 The regeneration was deferred until after EVA-3.  Metox
canisters were used in the JAL, and two fresh canisters were
delivered by the Shuttle.

Tomorrow’s EVA has the objective to replace the failed wrist roll
(WR) joint of the Canadian SSRMS (space station remote manipulator
system) with a new joint, delivered on STS-111.  The spacewalk
is to get underway at 10:30 am EDT and is expected to last seven
hours. Diagnostics software was used to perform a test on the WR
joint twice, and telemetry data continue to point toward a persistent
short in the joint.  The data will be used to troubleshoot the
returned joint.  The SSRMS is currently powered down, while the
normal OCS software is being restored.  It will subsequently be
re-powered and checked out to verify that the normal software is

Shuttle reserves indicate that we are getting close to being able to
add an extension day if required.

At 10:58 am EDT, members of the STS-111 flight crew participated in
an interactive televised PAO event with ABC Radio (Andy Field), WTOP
Radio/Washington, D.C. (Dimitri Sotis), and CNN/Atlanta (Miles
O’Brien).  The interview lasted 20 minutes, i.e., 6 min. per
media client.

Later in the day, at 3:53 pm EDT, all crew members are scheduled to
join in a 45-min. news conference with press media at NASA centers,
CNES/Paris and CSA/Canada.  

On Friday, 6/14, an American flag recovered from the debris of the
World Trade Center will be returned to New York City during a special
presentation scheduled for Flag Day.  The American Museum of
Natural History will host the event. The flag was flown into space in
December on STS-108, along with other commemorative badges, patches
and nearly 6,000 smaller American flags that will be presented to the
victims’ families in New York, Washington and Pennsylvania.  The
event, at 2:30 pm EDT, will be broadcast live on NASA TV.

Target areas of the Russian Uragan (Hurricane) earth observation
program on today’s task list of CDR Korzun were the
Bazaro-Karabulaksk region in Saratov province (strip photography),
Ilek-Aksai, and quarries near Leipzig and Dresden in Germany.

Today’s targets of the combined U.S. Shuttle/ISS earth observations
programs were Egypt, Toshka Lakes, Lake Nasser (in less
than three years several new lakes have been created in the desert
west of Lake Nasser as part of an ambitious resettlement and economic
program of the Egyptian government. Detailed photos of the flooded
areas and of any new construction activities [roads, canals,
settlements] along the shores are needed. Space photos will
complement our growing time series of this rapidly changing
, Caspian Sea, Kara Bogaz Gol (E. Caspian Sea
level has been rising during the past two decades, flooding low-lying
grazing and crop lands.  The Gulf of Kara Bogaz Gol was an
ephemeral salt pan until a channel was cut through the sand spit that
separates it from the Caspian. The channel was cut to allow rising
Caspian waters to spread into the Gulf.  Panoramic views of the
Gulf and detailed photos of the changing coastline will be used by
Russian colleagues investigating the rise of the Caspian)
Aral Sea (the rapidly disappearing Aral Sea lies left of
track, affording a chance for a panoramic view to the north. Detailed
views of the Amu Darya river delta will record any river waters that
may be reaching the sea. Winds commonly whip up the former seafloor
sediments, which are laden with agricultural and industrial chemical
waste [e.g., mercury], and blow them over croplands on the north
shore.  Of interest: recording any dust storms in the
, Ascension Island (this small island on the
Mid-Atlantic Ridge will be immediately left of track. Detailed photos
are needed of the volcanic island and the surrounding shoals).
Crete, Cyclades Islands, Turkish Coast (ISS track
followed the zone where the African tectonic plate is descending
beneath Eurasia, producing the islands of Crete, the Cyclades and the
Dodecanese off the Turkish coast. Crete and the volcanic island of
Santorini [Thira] were right of track; the Peloponnesos and Attic
Peninsula [Athens] were left. Offshore from Turkey will be Chios and
Lesvos. Most islands consist of seafloor sediments/lavas that have
been crumpled, uplifted above the sea. Detailed views of volcanoes,
faults are needed
), Balearic Islands, Corsica, N. Italy,
Po/Adriatic Smog
(the clouds that have blanketed Europe for
the past several days are moving out to the east, providing an
opportunity to photograph the Balearic Islands off Spain, Corsica and
northern Italy. The island of Elba, Florence and the diminutive
nation of San Marino were immediately left of track. Crew to watch
for any new smog accumulations in the Po River valley, which may
begin to drift down the Adriatic Sea. Slight overexposure will permit
identification of the feather edges of aerosol masses)
, Lake
Titicaca, Amazon Drainage Basin
(winter water levels in Lake
Titicaca should be documented. Clear weather prevails over the
southern headwaters of the Amazon  crew to look north for views
of megafans coming off the eastern Andes, then to follow Rio Madre de
Dios and Rio Madeira to the confluence with the mighty Amazon. As the
ISS approached the coast, the crew was to look right for views of the
vast Amazon delta. Devils Island were just left of track off the
, France, Brittany (the crew encountered the
coast of France near Nantes, with the Brittany coast on the left and
the Gironde estuary to the right. Clouds were moving out of the
region, permitting detailed nadir views of Nantes, Le Mans and Paris.
Detailed views of the upper Rhein rift, near the right-angle bend in
the Rhein west of Frankfurt, are needed — particularly the rarely
photographed volcanic rocks directly south of the river bend)
S. California Fires, E. Sierra Nevada Lakes (the crew
should have had near-nadir views of wildfires in the hills north and
east of Los Angeles — any good ESC [electronic still camera] images
that they downlink would be candidates for submission to the Earth
Observatory website. Mono Lake on the eastern flank of the Sierra
Nevada were just left of track. Crew to document water level and any
early summer vegetation in the vicinity)
, Hawaii, Kilauea
Volcano, Wildfires
(as the station passed over Maui, the Big
Island were to the right. Kilauea has sent new lava flows down to the
coast within the past week and steam plumes have been common.
Wildfires are also burning on the island. As with the LA wildfires,
good ESC images of the Hawaiian fires would be possible candidates
for Earth Observatory)
, Oregon Coast, Cascade Volcanoes
(the crew encountered the craggy Oregon coast near Portland, City
of Roses. The crew was to look both left and right of track and sight
along the line of volcanoes stretching to the north [St. Helens, Mt.
Adams, Mt. Rainier, Mt. Baker] and south [Mt. Hood, Mt. Jefferson].
This straight string of volcanoes rises along the continental margin
where the Juan de Fuca oceanic plate descends [is subducted] beneath
the North American tectonic plate. The Columbia River Gorge was right
of track as the ISS headed toward Spokane)
, S. Australia,
Nullarbor Escarpment, Ephemeral Lakes
(as  the crew
encountered the Australian coast at the Eyre Peninsula, they were to
look left along the sea cliffs of the ragged coast — the Nullarbor
Escarpment, which rises as much as 500 ft above the sea. This feature
is rarely photographed and long panoramic views are needed. Ephemeral
lakes Gairdner, Acraman [a meteor impact crater], Torrens and Frome
were left of track.  Crew to record any water in these lakebeds,
as well as the various shorelines. Lake Eyre was in the left
, and Great Barrier Reef (most of the Great
Barrier Reef was to the left of the ISS track as it approached the
coast near Mackay. The station passed over the area where the lagoon
begins to narrow northward.  Detailed mapping views of patch
reefs, islets and the crest of the reef escarpment are needed. Coral
bleaching has been reported in parts of the reef, in response to
warming waters)
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 6:15 am

Environmental Control and Life Support (ECLSS) and
Thermal Control (TCS):
Elektron O2 generator is powered On (16-amp mode). Vozdukh
CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump failed).
 U.S. CDRA CO2 scrubber is Off.  BMP Harmful Impurities
unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.

SM Working Compartment:  Pressure (mmHg) — 753,
temperature (deg C) — 27.4, ppO2 (mmHg) — 146.0, ppCO2 (mmHg) —
SM Transfer Compartment:  Pressure (mmHg) — 752,
temperature (deg C) — 19.9.
FGB Cabin:  Pressure (mmHg) — 752, temperature (deg C)
— 19.0.
Node:  Pressure (mmHg) — 740.71, temperature (deg C) —
22.4 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab:  Pressure (mmHg) — 742.59, temperature (deg C)
— 24.6, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) — 742.49,
temperature (deg C) — 23.1; shell heater temp (deg C) — 22.4, ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1:  Shell heater temp (deg C) — 22.5
PMA-2:  Shell heater temp (deg C) — 19.8

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not
available because MCA [major constituent analyzer] is failed and in
Extended Life mode
[= a state that preserves mass spectrometer
vacuum but produces no pp data]). MSA (mass spectrometer assembly)
and VGA (verification gas assembly) have been removed for return to

Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in
Autotrack mode.
SM batteries: (data not available)
FGB:  All batteries (6) are in “full charge” mode.
Plasma Contactor Unit PCU-1 in Discharge mode; PCU-2 in Discharge

Thermal Control Systems:
Air conditioner SKV-1 is Off. SKV-2 is On.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1
is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM on Standby.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:
All Russian communications & tracking systems are
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating

SSRMS/Canadarm2 positioned for Wrist Roll (WR) joint
RWS (robotics workstations) are On.
SSRMS WR joint to be replaced tomorrow EVA-3.

SpaceRef staff editor.